Switch outage in Kista Gate

Incident Report for SONIX

Resolved

This incident has been resolved.
Posted May 19, 2022 - 18:21 CEST

Update

New information: The fiber between Kista Gate and KN7 has suffered a cut and is expected to be repaired 02:00 CEST. Hopefully this will restore data plane service.
Posted May 16, 2022 - 21:42 CEST

Update

Today at 12:42 PM UTC (14:42 PM CEST) both links between Kista Gate and KN7 went down. Plan of action is being discussed.
Posted May 16, 2022 - 21:39 CEST

Investigating

We discovered that the management plane on the IXP switch located in Kista Gate has stopped sending logs and metrics to our collection daemons. We are investigating. At this time it seems the problem does not affect any data plane traffic.
Posted Apr 21, 2022 - 15:34 CEST
This incident affected: IX Platform (Obe Kista Gate).